### v1

# Baikal–GVD Collaboration Constraints on the diffuse flux of multi-PeV astrophysical neutrinos obtained with the Baikal Gigaton Volume Detector

Abstract
Various theoretical models predict cosmic neutrinos with multi-PeV energies. The recent detection of a eV neutrino in the KM3NeT experiment suggests that these energetic particles can be studied with present-day installations. Here, we present upper limits on the flux of astrophysical neutrinos with energies eV obtained with the largest liquid-water neutrino telescope, Baikal Gigaton Volume Detector (GVD), operation using cascade-like events. We discuss astrophysical implications of these results and constrain several cosmogenic neutrino scenarios using a combination of Baikal-GVD, KM3NeT, IceCube and Auger data.
I Introduction
Ultra-high energy (UHE) cosmic rays (CR) are accelerated in astrophysical sources and propagate towards the Earth from cosmological distances. Interactions of UHE particles with the Cosmic Microwave Background (CMB) radiation become important at high energies due to pion production for protons and spallation for heavier nuclei. For the uniform distribution of UHECR sources in the Universe, these interactions result in the suppression of the observed cosmic-ray flux above eV, known as the Greizen-Zatsepin-Kuzmin (GZK) cutoff [1, 2]. Such a suppression in the UHECR spectrum was observed by the High Resolution Fly’s Eye (HiRes) experiment [3], and later confirmed by the Pierre Auger Observatory (Auger) [4] and Telescope Array (TA) [5]. While the shape of the cutoff observed by HiRes and TA is consistent with the GZK prediction for the proton-dominated primary composition, the Auger observations favor the “disappointing scenario” [6] in which the suppression is caused by the maximal energy achieved in cosmic acceleration, and the composition is heavier.
These very same interactions produce secondary gamma rays and neutrinos from -meson decays. These cosmogenic neutrinos, pioneered by Berezinsky and Zatsepin [7], are now considered as an important diagnostic tool for UHECR models, see e.g. Ref. [8]. Indeed, the determination of the energy spectrum and the primary composition of highest-energy cosmic rays is plagued by systematic uncertainties related to the lack of firm understanding of hadronic interactions at energies beyond the reach of collider experiments, cf. e.g. Refs. [9, 10, 11]. The cosmogenic neutrino flux, if observed, could be used to infer cosmic-ray composition and source parameters in an independent way. Roughly, higher fluxes of cosmogenic neutrinos are associated with lighter composition and/or excess of sources at high redshifts, but the actual fluxes depend on many parameters, see e.g. Refs. [12, 13] and references therein. Cosmogenic neutrinos are expected to come from all UHECRs in the Universe, providing truly diffuse flux of neutrinos, as opposite to neutrinos from individual astrophysical sources.
Recently, the KM3Net experiment has reported about the detection of a neutrino with best-fit energy well above 100 PeV [14]. This event could potentially be the first cosmogenic neutrino ever detected [15]. However, to make conclusions from this observation, one should take into account that events of similar energies have not been detected by other neutrino telescopes, despite larger exposure. These non-detections constrain strongly the isotropic neutrino flux, see e.g. [16, 17].
Still under construction, Baikal-GVD [18] is the largest operating neutrino telescope in the Northern hemisphere, and has already collected significant exposure for high-energy neutrinos. No event with energy above eV has been detected by Baikal-GVD in the cascade mode so far, so we present here corresponding upper limits on the isotropic flux of neutrinos at highest energies, and discuss possible implications of this non-detection.
II Data
Baikal-GVD is a water Cerenkov neutrino telescope, currently having the instrumented volume km3 and growing. It is located in Lake Baikal, Russia, at the latitude of North.
For the present analysis, we use the Baikal-GVD sample of cascade events detected between Spring 2018 and Spring 2024. In this period, the detector was operated in several, increasingly larger configurations, with one or two new “clusters” of optical modules (OMs) added each spring. One cluster comprises 288 OMs arranged on eight vertical strings, with 36 OMs per string. Over the stated period, the number of clusters increased from three to eleven. The total exposure of the data set used here corresponds to years of one-cluster operation. The reader may find descriptions of the experiment, event reconstruction, analysis techniques and procedures e.g. in Refs. [18, 19, 20, 21, 22].
We select high-energy cascade events with the multiplicity of triggered optical modules (OMs) 8 and impose the quality cuts described in Ref. [19], except for the cut introduced to suppress atmospheric muon bundles, which are not expected to contaminate the sample at these high energies. Relaxing this latter cut and multiplicity of OMs have allowed to increase the exposure considerably.
The effective area of Baikal-GVD as of 2023, determined by Monte-Carlo simulations following Ref. [19], is presented in Fig. 1.
Note that for cascade events produced by electron (anti)neutrinos, the Landau-Pomeranchuk-Migdal (LPM) effect suppresses the effective area at high energies. The LPM effect was taken into account in simulations following Refs. [23, 24, 25]. For water, it starts to suppress bremsstrahlung for electron energies above PeV and therefore significantly reduces the effective area for electron (anti)neutrinos with energies above PeV. The peak of the effective area about 6.3 PeV is related to the resonant production of bosons by electron antineutrinos, known as the Glashow resonance.
III Flux limits
The highest-energy event found in the Baikal-GVD cascade sample had the energy of TeV; this event has already been reported in Ref. [19]. There are no events in the range of interest, TeV. The expected atmospheric background is negligible, so we use the standard approach for the Poisson statistics, recommended by Ref. [26], to obtain 90% CL upper limit on the Poisson mean of the observed number of events as . As it is customary in high-energy neutrino astrophysics, we derive our limit for overlapping decade-wide energy bins, assuming spectrum within each bin to account for the energy dependence of the exposure, that is
where , is the effective area, the integration extends over the bins with in energy and over the data collection time in . The resulting upper limits are listed in Table 1
| energy bin, | upper limit, |
| TeV | GeV cm-2s-1sr-1 |
| 0.78 | |
| 1.6 | |
| 2.5 | |
| 4.3 | |
| 7.6 | |
| 14 | |
| 27 | |
| 54 |
and presented graphically in Fig. 2.
IV Discussion
IV.1 Baikal-GVD limits and results of other experiments
Baikal-GVD upper limits on the isotropic neutrino flux are compared in Fig. 2 with the observation by KM3NeT [14] and latest upper limits from IceCube [28] and Auger [27]. Baikal-GVD limits are becoming stronger as compared to Auger limits at the energies below 100 PeV. Our upper limit is comparable to that of IceCube in the lowest energy bin, and is not in tension with the IceCube observation of several events at those energies [28].
Following Ref. [16], we estimate the neutrino flux at the energies between 72 PeV and 2600 PeV, that is in the 90% CL energy interval of KM3-230213A. We assume that this event originates from the isotropic, steady flux of neutrinos equally distributed between flavors. The energy-dependent effective areas for KM3NeT [14] (335 days exposure), IceCube [28] (4605 days), Auger [27] (18 years), and Baikal-GVD (this work; 26.8 cluster-years) are presented in Fig. 3.
Integrating the total exposure over energy, assuming neutrino flux, and having one observed event for zero background, we obtain the estimate
It may be compared with the KM3NeT-only estimate in Fig. 4.
IV.2 Constraints on the cosmogenic neutrino flux models
Predictions of the cosmogenic neutrino flux depend strongly on the assumed properties (injection spectra and composition) and spatial distribution (cosmological evolution and number density) of UHECR sources, as well as on the chosen model of the cosmic photon background, see e.g. [12]. From a plethora of available theoretical predictions, we have chosen five recent models to compare with the new experimental constraints.
Auger 2023 low and high.
Ref. [29] presents a number of cosmogenic neutrino fluxes derived from simultaneous fits to spectra and composition observed by Auger, assuming two components from different populations of sources with varying parameters. These predictions span a wide band between the lowest and highest fluxes, both corresponding to the strong positive evolution, , of the low-energy component with the redshift , and the flat evolution of the high-energy component, but differing by assumed maximal source redshifts, for the low-flux model versus for the high-flux one.
Muzio et al. 2023 maximal.
Ref. [30] considers a particular model in which cosmic rays interact efficiently near their sources, and assume a significant injected proton component at super-GZK energies. We use the model providing for the maximal neutrino flux, which corresponds to the AGN evolution and EPOS-LHC interaction model.
Ehlert et al. 2023 2SC-dip.
Ref. [13] performs a comprehensive study of possibilities to have a significant UHECR proton component without contradicting to observational data. Interestingly, their UHECR best fit within one of the two models considered, 2SC-dip, is in tension with observational constraints on the neutrino flux. We use this best-fit model prediction without multmessenger constraints because they are being reconsidered in the present study.
KM3NeT 2025 extreme evolution.
Ref. [15] extended the source-evolution parameter space in the Auger data fit to obtain the extreme scenario with strong positive evolution and , in which cosmogenic neutrinos could explain the KM3-230213A event, which we also consider.
Predictions of the selected models are shown in gray color in both Fig. 2 and Fig. 4. To obtain quantitative constraints on the scenarios, we calculate, for each of them, the expected number of events in the KM3-230213A 90% CL energy uncertainty band. Table 2
| best fit for | upper limit | |
| model | cosmogenic | (90% CL) |
| KM3-230213A | ||
| Auger low 2023 [29] | 610 | |
| Auger high 2023 [29] | 11.5 | |
| Muzio maximal 2023 [30] | 0.56 | |
| Ehlert dip 2023 [13] | 0.96 | |
| KM3NeT 2025 extreme [15] | 3.9 |
presents the estimates of model normalization factors obtained assuming that the flux corresponding to a single detected event, KM3-230213A, is cosmogenic. Within this assumption, the lowest-flux Auger 2023 model is disfavored at . We also present there upper limits on the normalization factors, which are free from this assumption.
V Conclusions
This work presents upper limits on the isotropic astrophysical neutrino flux at energies above TeV from searches of neutrino-induced cascades in the 2018–2023 data of Baikal-GVD in the construction stage, equivalent to years of the present configuration of the detector. In the lower part of the energy band, these limits are of the same order as those reported by IceCube, and supersede those published by Auger. We further concentrate on the energy range in which the KM3-230213A event was detected. We combine the exposure of Baikal-GVD with those of IceCube, Auger, and KM3NeT, and present the best-fit isotropic flux estimate from the joint analysis. With the combined exposure, we also obtain constraints on selected models of cosmogenic neutrinos; other models for the isotropic neutrino flux can be tested in the same way with the help of the presented energy-dependent effective areas. Future analyses are needed to understand the origin of KM3-230213A and to obtain more information on UHECR sources with the neutrino data.
Acknowledgements
This work is supported in the framework of the State project “Science” by the Ministry of Science and Higher Education of the Russian Federation under the contract 075-15-2024-541. The work of A.V.P. was supported by the Black Hole Initiative, which is funded by grants from the John Templeton Foundation (Grant #60477, 61479, 62286) and the Gordon and Betty Moore Foundation (Grant GBMF-8273).

### v2

@@
 Cosmogenic neutrinos are expected to come from all UHECRs in the Universe, providing truly diffuse flux of neutrinos, as opposite to neutrinos from individual astrophysical sources.
+However, this cosmogenic neutrino flux is yet to be measured.
+Only three neutrino events with best-fit reconstructed energies above 5 PeV have been detected by the presently largest neutrino telescope, IceCube.
+The highest energy of these three is about 13 PeV, see e.g.
+Ref.
+[14] for details and further references.
 Recently, the KM3Net experiment has reported about the detection of a neutrino with best-fit energy well above 100 PeV [15].
@@
 One cluster comprises 288 OMs arranged on eight vertical strings, with 36 OMs per string.
-Over the stated period, the number of clusters increased from three to eleven.
+The layout of the detector is presented in Fig.
+1.
+The inter-cluster spacing is varied in the range 250–300 m.
+Given the typical geometrical size of the cascade, the installation effectively works as a set of independent clusters in the cascade mode.
+Over the stated period, the number of clusters increased from three to twelve.
 The total exposure of the data set used here corresponds to years of one-cluster operation.
@@
 [19, 20, 21, 22, 23].
-We select high-energy cascade events with the multiplicity of triggered optical modules (OMs) 8 and impose the quality cuts described in Ref.
-[19], except for the cut introduced to suppress atmospheric muon bundles, which are not expected to contaminate the sample at these high energies.
-Relaxing this latter cut and multiplicity of OMs have allowed to increase the exposure considerably.
+The search strategy for high-energy neutrino detection with Baikal-GVD was described in Ref.
+[20].
+High-energy cascade events with the multiplicity of triggered optical modules (OMs) 19 and reconstructed energy TeV were selected, and additional cuts which suppress events from atmospheric muons were applied.
+The cut 19 was optimised for effective suppression of misreconstructed background events from atmospheric muon bundles.
+The number of these background events decrees rapidly with energy due to the steep energy spectrum of atmospheric muons.
+In the present work, we select high-energy cascade events with 8 and impose the quality cuts described in Ref.
+[20], except for the cut introduced to suppress atmospheric muon bundles.
+Relaxing this latter cut and multiplicity of OMs allows us to increase the exposure considerably.
+Simulations indicate that the number of background events is about for all energies above 3 PeV, and above 10 PeV with these relaxed cuts.
 The effective area of Baikal-GVD as of 2023, determined by Monte-Carlo simulations following Ref. [20], is presented in Fig. 2.
@@
 As it is customary in high-energy neutrino astrophysics, we derive our limit for overlapping decade-wide energy bins, assuming spectrum within each bin to account for the energy dependence of the exposure, that is
-where , is the effective area, the integration extends over the bins with in energy and over the data collection time in . The resulting upper limits are listed in Table 1
+where , is the effective area, the integration extends over the bins with in energy, over the celestial sphere in the solid angle and over the data collection time in .
+The effective area, measured in m2, is estimated by Monte-Carlo simulations and takes into account the efficiency of neutrino detection for a particular set of selection cuts.
+The resulting upper limits are listed in Table 1
 | energy bin, | upper limit, |
@@
 The energy-dependent effective areas for KM3NeT [15] (335 days exposure), IceCube [29] (4605 days), Auger [28] (18 years), and Baikal-GVD (this work; 26.8 cluster-years) are presented in Fig.
-3.
+4.
 Integrating the total exposure over energy, assuming neutrino flux, and having one observed event for zero background, we obtain the estimate
@@
 3 and Fig.
-4.
-To obtain quantitative constraints on the scenarios, we calculate, for each of them, the expected number of events in the KM3-230213A 90% CL energy uncertainty band.
-Table 2
-| best fit for | upper limit | |
-| model | cosmogenic | (90% CL) |
-| KM3-230213A | ||
-| Auger low 2023 [29] | 610 | |
-| Auger high 2023 [29] | 11.5 | |
-| Muzio maximal 2023 [30] | 0.56 | |
-| Ehlert dip 2023 [13] | 0.96 | |
-| KM3NeT 2025 extreme [15] | 3.9 |
-presents the estimates of model normalization factors obtained assuming that the flux corresponding to a single detected event, KM3-230213A, is cosmogenic.
-Within this assumption, the lowest-flux Auger 2023 model is disfavored at .
-We also present there upper limits on the normalization factors, which are free from this assumption.
+5.
+To obtain quantitative constraints on the scenarios, we make use of Eq.
+(1) and substitute there the theoretically predicted flux .
+In this way, we calculate, for each of the models, the expected number of events in the KM3-230213A 90% CL energy uncertainty band, in which the observed number of events is one.
+These theoretical numbers are presented in Table 2.
+| model | |||
+| Auger low 2023 [30] | 0.006 | 610 | |
+| Auger high 2023 [30] | 0.34 | 11.5 | |
+| Muzio maximal 2023 [31] | 7.0 | 0.56 | |
+| Ehlert dip 2023 [13] | 4.1 | 0.96 | |
+| KM3NeT 2025 extreme [16] | 0.98 | 3.9 |
+Assuming that the model cosmogenic neutrino spectrum scales with the normalization factor , where for the theoretical model and the spectral shape does not change, we obtain 90% CL upper limits on from Eq.
+(1).
+We then make a temporary assumption that the single observed event was cosmogenic, and use the Poisson statistics [27] to get the best-fit value of and its 68% confidence interval.
+These numbers are also presented in Table 2.
+We see that the assumption of the cosmogenic origin of KM3-230213A is in tension with the lowest-flux Auger 2023 model, with the Poisson p-value corresponding to about discrepancy.
 V Conclusions
@@
 was supported by the Black Hole Initiative, which is funded by grants from the John Templeton Foundation (Grant #60477, 61479, 62286) and the Gordon and Betty Moore Foundation (Grant GBMF-8273).
+Data availability
+The effective area of the Baikal-GVD experiment is available in an ancillary file with this paper’s arXiv version [33].